in a survey , 65% of Americans said they own an answering machine. If 20 Americans are selected randomly a) find the probabilities that at least 18 people own an answering machine and b) find the mean and the standard deviation the number of people owning an answering machine .
Answer by stanbon(57307)

in a survey , 65% of Americans said they own an answering machine.
-----------------------------------------------
If 20 Americans are selected randomly
a) find the probabilities that at least 18 people own an answering machine
P(18<= x <=20) = 1-binomcdf(20,0.65,17) = 0.0121
------------------------------------------------------
b) find the mean and the standard deviation the number of people owning an answering machine .
mean = np = 20*0.65 = 13
std = sqrt(13*0.35) = 2.1331
===============================
cheers,
Stan H.
------------------------------